Calculate the third Bohr's energy of the hydrogen atom
1
Expert's answer
2020-05-18T10:14:08-0400

The orbital energy for nth level is


"E_n=-\frac{2\pi e_e^4m_ek^2}{h^2n^2}=-\frac{13.6\text{ eV}}{n^2},\\
\space\\
E_3=-\frac{13.6}{3^2}=-1.51\text{ eV}."